# Spatially covariant gravity with two degrees of freedom in the presence of an auxiliary scalar field: Hamiltonian analysis

### v1

@@ I Introduction
With the detection of gravitational waves, a novel tool has emerged for testing theories of gravity. A fundamental question arises: Is general relativity the only theory of gravity that propagates gravitational waves with precisely two degrees of freedom (2DOFs)? This question is partially addressed by Lovelock’s theorem [1, 2], which states that in four dimensions, if one requires spacetime covariance and second-order equations of motion, general relativity is indeed the unique theory that propagates gravitational waves with two degrees of freedom. This also implies that, by violating the assumptions of Lovelock’s theorem, other gravitational theories propagating two degrees of freedom can, in principle, exist.
Gravitational theories with 2DOFs can be traced back to the Cuscuton theory [3], which has been extensively studied [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16], and extended in the framework of scalar-tensor theories with higher-order derivatives [17, 18]. Such theories have also been discussed as special cases of Hořava gravity [19, 20, 21, 22]. The connection between 2DOFs and spacetime symmetries has been investigated in [23, 24, 25, 26].
Another class of theories propagating 2DOFs was proposed in [27], known as minimally modified gravity (MMG) [^1] See also the minimal theory of massive gravity studied in [28, 29, 30, 31, 32, 33].. The so-called type-II MMG [34, 35] includes the original Cuscuton theory as a special case. MMG has been extensively studied [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51], and has also been generalized by introducing auxiliary constraints in phase space [52, 53].
@@
A general framework for spatially covariant gravity (SCG), which respects only spatial covariance, was proposed in [54, 55]. Due to the breaking of full spacetime covariance, SCG generally propagates a scalar degree of freedom in addition to the 2DOFs corresponding to gravitational waves. Thus, it provides an alternative approach to constructing scalar-tensor theories [56, 57, 58, 59]. SCG includes Hořava gravity [60, 61, 62] and the effective field theory of inflation/dark energy [63, 64, 65, 66, 67] as special cases. It has been further generalized to include a dynamical lapse function [68, 69], nonmetricity [70], parity violation [71], and multiple scalar modes [72]. SCG has also been applied in cosmology and gravitational wave studies [73, 74, 75, 76, 77]. Subclasses of SCG that propagate only 2DOFs were explored in [78, 79], and in the presence of a dynamical lapse function in [80]. The explicit Lagrangian found in [78] has been employed in cosmological studies [81, 82, 83, 10, 84].
@@
In this work, we generalize the SCG framework to include an auxiliary scalar field and determine the conditions under which the theory propagates only 2DOFs. The idea of spatially covariant gravity with an auxiliary scalar field was first introduced in [85], originally motivated by generally covariant scalar-tensor theories in which the scalar field has a spacelike gradient [^2] However, we emphasize that these are entirely different theories. See Appendix A for details.. The presence of an auxiliary scalar field introduces new and novel possibilities for constructing 2DOF theories.
@@
In [86], a perturbative analysis similar to [17, 69, 79] was performed to derive necessary conditions for the absence of scalar modes at linear order in perturbations around a cosmological background. Nevertheless, a Hamiltonian constraint analysis is still required to determine the necessary and sufficient conditions for the theory to propagate only 2DOFs at the nonlinear level. This paper is devoted to this task.
@@
The paper is organized as follows. In Sec. II, we describe our model of spatially covariant gravity with an auxiliary scalar field. We find that new terms in the Lagrangian are allowed due to the inclusion of the auxiliary scalar. In Sec. III, we derive the Hamiltonian formalism and show that, in general, a scalar mode propagates unless additional constraints are imposed. In Sec. IV, we perform a degeneracy analysis and derive the conditions under which the theory propagates only 2DOFs. In Sec. V, we apply our formalism to a concrete model, where the Lagrangian is built from SCG monomials with $d=2$. Finally, in Sec. VI, we summarize our results.
@@ II Spatially covariant gravity with an auxiliary scalar field
where $N$ is the usual lapse function, $h_{ij}$ is the spatial metric on the spacelike hypersurfaces, ${}^{3}\!R_{ij}$ is the spatial Ricci tensor, $\mathrm{D}_{i}$ is the spatially covariant derivative compatible with $h_{ij}$, and £$\pounds_{\bm{n}}$ is the Lie derivative with respect to the normal vector $n^{a}$ of the hypersurfaces. By definition, the action (1) respects only the three-dimensional spatial invariance. We require that the Lie derivative £$\pounds_{\bm{n}}$ acts only on $h_{ij}$, which yields the extrinsic curvature $K_{ij}$ defined by
£$\pounds_{\bm{n}}h_{ij}=2K_{ij}=\frac{1}{N}\left(\dot{h}_{ij}-\pounds_{\vec{N}}h_{ij}\right).$ (2)
As a result, both the lapse function $N$ and the scalar field $\phi$ play the role of auxiliary variables. Note that the shift vector $N^{i}$ should not appear explicitly in the Lagrangian, as it is actually the gauge field of spatial covariance.
Since the Lie derivative enters the Lagrangian only through the extrinsic curvature, we simply introduce $K_{ij}$ to replace the Lie derivative £$\pounds_{\bm{n}}$ in (1), and consider the following action:
@@
According to the Hamiltonian analysis in the next section, the action (3) generally has 3 DOFs. This is simply because $\phi$ has no dynamics; the number of DOFs in (3) is the same as in spatially covariant gravity without the scalar field [54, 55].
In principle, Lie derivatives of the lapse function £$\pounds_{\bm{n}}N$ and of the extrinsic curvature £$\pounds_{\bm{n}}K_{ij}$ can be considered, which would introduce additional degrees of freedom. It has been shown that extra conditions must be imposed in order to eliminate these unwanted degrees of freedom. For example, the theory will propagate an extra DOF (i.e., a total of 4 DOFs) in the case of a dynamical lapse function. Two conditions are required to eliminate such an extra DOF [68]. A similar analysis was carried out to fully eliminate the scalar DOFs [80].
Generally, the Lagrangian in (3) may be nonlinear in $K_{ij}$, which makes the explicit inversion of velocity $\dot{h}_{ij}$ in terms of the momentum $\pi^{ij}$ impossible. This problem can be solved by rewriting (3) in an equivalent form:
@@
where the action $S\left(\phi,N,h_{ij},{}^{3}R_{ij},B_{ij};\mathrm{D}_{i}\right)$ is the same as (3) with $K_{ij}$ replaced by $B_{ij}$. This strategy of using a Lagrange multiplier to linearize the velocities was employed in [68, 78]. The equations of motion for $\varLambda^{ij}$ ensure that the auxiliary field $B_{ij}$ is exactly equal to $K_{ij}$. Thus, the two actions are equivalent, at least at the classical level. By making use of the equation of motion for the auxiliary field $B_{ij}$, we can determine the Lagrange multiplier $\varLambda^{ij}$ as
@@
Finally, by plugging the solution for $\varLambda^{ij}$ into the action (4), we can recast it in the form
@@ III.1 Hamiltonian and primary constraints
and the corresponding set of conjugate momenta is denoted by
@@
Since all velocities $\dot{\varPhi}^{I}$ in the Lagrangian (6) appear linearly, the conjugate momenta automatically yield constraints among the phase space variables. Indeed, the conjugate momenta are
@@
which lead to the primary constraints
@@
Here and throughout this work, the symbol $\approx$” denotes “weak equality”, i.e., an equality that holds only when the constraints are satisfied. By contrast, “$=$” denotes strong equality, which holds regardless of whether constraints are imposed.
@@
and define the subspace of the phase space where the primary constraints are satisfied as $\varGamma_{\mathrm{P}}$, which can be regarded as a hypersurface in the original phase space $\left\{\varPhi^{I},\varPi_{I}\right\}$.
The canonical Hamiltonian is obtained via the Legendre transformation:
@@
$=$ $\int\mathrm{d}^{3}x\left(\pi_{i}\pounds_{\vec{\xi}}N^{i}+p\pounds_{\vec{\xi}}\phi+p^{ij}\pounds_{\vec{\xi}}B_{ij}+\pi\pounds_{\vec{\xi}}N+\pi^{ij}\pounds_{\vec{\xi}}h_{ij}\right).$
In the following, we briefly summarize the properties of $X[\vec{\xi}]$ as discussed in [68]. Suppose $Q\left(\vec{x}\right)$ is a general spatial tensor field or density (with spatial indices suppressed for brevity) constructed from the phase space variables $\left\{\varPhi^{I},\varPi_{I}\right\}$, and let $\mathcal{F}$ be a scalar functional defined by
@@
where $f(\vec{x})$ is a smooth test function. Note that $f(\vec{x})$ may also carry appropriate spatial indices (suppressed here) such that $\mathcal{F}$ is invariant under time-independent spatial diffeomorphisms. The Poisson bracket between $X[\vec{\xi}]$ and $\mathcal{F}$ is given by
@@
In (18), the Poisson bracket $[\mathcal{F},\mathcal{G}]$ is defined as
$[\mathcal{F},\mathcal{G}]\coloneqq\int\mathrm{d}^{3}z\left(\frac{\delta\mathcal{F}}{\delta\varPhi_{I}(\vec{z})}\frac{\delta\mathcal{G}}{\delta\varPi^{I}(\vec{z})}-\frac{\delta\mathcal{F}}{\delta\varPi^{I}(\vec{z})}\frac{\delta\mathcal{G}}{\delta\varPhi_{I}(\vec{z})}\right).$ (19)
A key point is that the right-hand side of the second line in (18) vanishes on the constraint surface if $Q(\vec{x})$ is a constraint.
By integrating by parts, $X[\vec{\xi}]$ can be written in a more convenient form:
@@
As we shall see later, the constraints $\mathcal{C}_{i}\approx 0$ are secondary constraints associated with the primary constraints $\pi_{i}\approx 0$. Furthermore, the Poisson bracket between $\mathcal{C}_{i}$ and any constraint $Q$ is given by
@@
Due to the presence of primary constraints, the time evolution is governed by the total Hamiltonian, defined as
@@
where $\lambda_{I}:=\left\{v^{i},v,v_{ij},\lambda,\lambda_{ij}\right\}$ are undetermined Lagrange multipliers. Since the canonical Hamiltonian $H_{\mathrm{C}}$ is only well-defined on the subspace defined by the primary constraints $\varGamma_{\mathrm{P}}$, we may directly use $\left.H_{\mathrm{C}}\right|_{\varGamma_{\mathrm{P}}}$ given in (14) in place of $H_{\mathrm{C}}$ in (13) for subsequent calculations.
@@ III.2 Consistency condition
Constraints must be preserved under time evolution. For the primary constraints, we require that
@@
which are the so-called consistency conditions for the primary constraints. These conditions may yield further constraints.
Without imposing additional assumptions on the structure of the Lagrangian in (6), the Poisson brackets between the primary constraints have been computed in detail in [85], and are summarized in Appendix B. For later convenience, we rewrite (25) in matrix form:
@@
where $\mathcal{C}_{i}$ is defined in (22). Note that $\mathcal{C}_{i}$ must have vanishing Poisson brackets with all other constraints due to the property (22). In the fourth line of (26), $\mathcal{C}$ is defined by
@@
According to (129), $\left[p^{ij}(\vec{x}),\tilde{\pi}^{kl}(\vec{y})\right]$ in the third line of (26) is proportional to the kinetic matrix
@@
which we assume to be non-degenerate. This is because $B_{ij}$ is the auxiliary field corresponding to $K_{ij}$. The non-degeneracy of (29) implies that the kinetic term for $\dot{h}_{ij}$ is non-degenerate, which guarantees that the theory includes General Relativity as a limiting case, and in particular, allows the existence of gravitational waves.
As a result, the Lagrange multipliers $\lambda_{ij}$ can be fixed from the third line of (26) as
@@
Then the second and fourth lines of (26) yield two additional secondary constraints:
@@
The constraints $\frac{\delta S}{\delta\phi(\vec{x})}$ and $\mathcal{C}(\vec{x})$ do not generate further constraints via consistency conditions, as will become clear in the next section.
Having identified all the primary and secondary constraints, we are now in a position to classify them. According to the terminology of the Dirac-Bergmann algorithm, a constraint that has vanishing Poisson brackets with all other constraints (including itself) is called a first-class constraint; otherwise, it is second-class. The distinction is that each first-class constraint eliminates one degree of freedom, whereas a second-class constraint removes one phase space dimension, i.e., half a degree of freedom.
From the discussion above on $\mathcal{C}_{i}$ and the constraint algebra in Appendix B, the Poisson brackets among all the constraints can be summarized in the so-called Dirac matrix:
@@
where an “$X$” denotes a generically non-vanishing Poisson bracket. In total, there are 22 constraints, which can be divided into two classes:
@@
The number of degrees of freedom (DOF) is thus
#$\#_{\mathrm{DOF}}$ $=$ $\frac{1}{2}\left(2\times\#_{\mathrm{var}}-2\times\#_{1\mathrm{st}}-\#_{2\mathrm{nd}}\right)$ (33)
@@
$=$ 3 3,3 ,
which explicitly shows that a spatially covariant gravity theory where temporal derivatives arise only through the extrinsic curvature has 3 propagating degrees of freedom, regardless of the presence of additional auxiliary field(s).

### v2

@@ I Introduction
In light of the detection of gravitational waves, gravitational waves provide a novel tool for testing theories of gravity. General relativity possesses only two propagating degrees of freedom (DOFs), which manifests as tensorial gravitational waves (GWs) in a homogeneous and isotropic background. The detection of GWs raises a fundamental question: Is general relativity the only theory of gravity that propagates GWs with precisely two degrees of freedom (2DOFs)? This question is partially answered by Lovelock’s theorem [1, 2], which states that in four dimensions, requiring spacetime covariance and second-order equations of motion, general relativity is indeed the unique theory propagating gravitational waves with two degrees of freedom. This also implies that, by violating the assumptions of Lovelock’s theorem, other theories of gravity propagating 2DOFs can, in principle, exist. From the theoretical point of view, a Lorentz covariant massless spin-2 particle possesses two physical polarization states. If one constructs a Lorentz invariant theory for such a massless spin-2 particle and further demands that it can be consistently coupled to matter fields, the resulting structure is necessarily that of general relativity. In this sense, general relativity as a theory of gravity is unique since its form is inextricably related to the masslessness of the graviton and the Lorentz symmetry of spacetime. Consequently, alternative theories of gravity that also propagate only 2DOFs provide a valuable testing ground for the foundation principles of general relativity, such as Lorentz invariance, diffeomorphism invariance, and the masslessness of graviton, etc.
Gravitational theories propagating 2DOFs can be traced back to the Cuscuton theory [3], which is introduced as a scalar field model with a non-canonical kinetic term $\propto\sqrt{\left|-\partial_{\mu}\phi\partial^{\mu}\phi\right|}$. The cuscuton theory represents the incompressible (infinite speed of sound) limit of $k$-essence, in which the scalar perturbation propagates with an infinite speed of sound in a cosmological background. Although linear perturbations exhibit superluminal propagation, the local phase space degenerates, implying no independent dynamical DOFs and thus preserving causality. As a result, the cuscuton theory modifies gravitational dynamics without introducing additional local DOFs. The Cuscuton theory was extensively studied [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16] and was extended in the framework of scalar-tensor theory with higher order derivatives [17, 18]. Such kind of theory has also been discussed as a special case of Hořava gravity [19, 20, 21, 22]. The relation between the 2DOFs and the spacetime symmetry has been investigated in [23, 24, 25, 26]. Another class of of theories propagating 2DOFs was proposed in [27], which is dubbed the minimally modified gravity (MMG) [^1] See also the minimal theory of massive gravity studied in [28, 29, 30, 31, 32, 33].. As the name suggests, MMG theories modify generally relativity “minimally” in the sense that they modify the gravitational sector without introducing extra DOFs beyond the standard 2DOFs. MMG theories extend general relativity by breaking time diffeomorphism invariance while preserving spatial diffeomorphism invariance, and whose action is linear in the lapse function. A key result is the derivation of a self-consistency condition, which ensures the theory possesses no more than two local physical DOFs. Examples of MMG include general relativity, square-root gravity, exponential gravity, and theories with lapse-independent terms, all of which exhibit a constrained phase space structure that eliminates the scalar graviton through first- or second-class constraints. The so-called type-II MMG [34, 35] includes the original Cuscuton theory as a special case. The MMG has been studied extensively [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51]. It has also been generalized in [52, 53] by introducing auxiliary constraints in the phase space.
@@
A general framework of spatially covariant gravity (SCG) theories respecting only the spatial covariance was proposed in [54, 55]. Due to the violation of general covariance, SCG generally propagates a scalar degree of freedom (besides the 2DOFs corresponding to the gravitational waves) and thus can be viewed as an alternative approach to constructing the scalar-tensor theories [56, 57, 58, 59]. The SCG includes the Horava gravity [60, 61, 62], effective field theory of inflation/dark energy [63, 64, 65, 66, 67] as special cases. It was further generalized with a dynamical lapse function [68, 69], with nonmetricity [70], with parity violation [71] and with multiple scalar modes [72]. The SCG has been applied in the study of cosmology and gravitational waves [73, 74, 75, 76, 77]. Subclasses of SCG propagating only 2DOFs were explored in [78, 79] and with a dynamical lapse function in [80]. The concrete Lagrangian found in [78] has been applied in the study of cosmology [81, 82, 83, 10, 84].
@@
In this work, we will generalize the SCG in the presence of an auxiliary scalar field, and determine the conditions such that the theory propagates only 2DOFs. The idea of spatially covariant gravity with an auxiliary scalar field was firstly introduced in [85], which was originally motivated by the generally covariant scalar-tensor theory when the scalar field possess a spacelike gradient [^2] However, we would like to point out that they are completely different theories. Please refer to Appendix A for details. . The presence of an auxiliary scalar field gives us more and novel possibilities to build the theory propagating two degrees of freedom.
@@
In [86], by performing a perturbative analysis similar to [17, 69, 79], necessary conditions such that no scalar mode propagates at linear order in perturbations around a cosmological background were determined. Nevertheless a Hamiltonian constraint analysis is still needed in order to determine the necessary and sufficient conditions for the theory to propagate only 2DOFs at the nonlinear level. This paper is thus devoted to this issue.
@@
This paper is organized as follows. In Sec. II, we describe our model of spatially covariant gravity with an auxiliary scalar field. It was found that new terms in the Lagrangian are allowed thanks to the introduction of the auxiliary scalar field. In Sec. III, we derive the Hamiltonian formalism and show that the theory generally propagates a scalar mode if no further constraint is imposed. In Sec. IV, we make the degeneracy analysis and derive the conditions such that the theory propagate only 2DOFs. In Sec. V, we apply our formal analysis to a concrete model, of which the Lagrangian is built of SCG monomials of $d=2$. In Sec. VI, we summarize our results.
@@ II Spatially covariant gravity with an auxiliary scalar field
where $N$ is the usual lapse function, $h_{ij}$ is the spatial metric on the spacelike hypersurfaces, $^{3}\!R_{ij}$ is the spatial Ricci tensor, $\mathrm{D}_{i}$ is the spatially covariant derivative compatible with $h_{ij}$, and $\pounds_{\bm{n}}$ is the Lie derivative with respect to the normal vector $n^{a}$ of the hypersurfaces. By definition, the action (1) respects only the 3-dimensional spatial invariance. We require that the Lie derivative $\pounds_{\bm{n}}$ acts only on $h_{ij}$, which yields the extrinsic curvature $K_{ij}$ defined by
$\pounds_{\bm{n}}h_{ij}=2K_{ij}=\frac{1}{N}\left(\dot{h}_{ij}-\pounds_{\vec{N}}h_{ij}\right).$ (2)
As a result, both the lapse function $N$ and the scalar field $\phi$ play the role of an auxiliary variables. Note the shift vector $N^{i}$ should not appear explicitly in the Lagrangian, which is actually the gauge field of the spatial covariance.
In this work, we restrict our attention to the first-order Lie derivative to prevent higher-order time derivatives from appearing in the equations of motion. Moreover, since the Lie derivative is assumed to act solely on $h_{ij}$, it enters the Lagrangian exclusively through the extrinsic curvature (2). Therefore, we simply introduce $K_{ij}$ to replace the Lie derivative $\pounds_{\bm{n}}$ in (1), and consider the following action
@@
According to the Hamiltonian analysis in the next section, generally the action (3) has 3 DOFs. This is simply because $\phi$ has no dynamics, the number of DOFs of (3) is the same as the spatially covariant gravity without the scalar field [54, 55]. In particular, since the extrinsic curvature is linear in the Lie derivative (and thus in time derivative), the resulting equations of motion from the action (3) contain at most second-order time derivatives. As a result, the theory avoids the Ostrogradsky ghost instability.
In principle, Lie derivatives of the lapse function $\pounds_{\bm{n}}N$ and the extrinsic curvature $\pounds_{\bm{n}}K_{ij}$ can be considered, which will introduced more degrees of freedom. It was shown that extra conditions must be put in order to evade these unwanted degrees of freedom. For example, the theory will propagate an extra DOF (i.e., in total 4 DOFs) in the case of a dynamical lapse function. Two conditions are needed in order to evade such an extra DOF [68]. Similar analysis was considered in order to fully eliminate the scalar DOFs [80].
Generally, the Lagrangian in (3) may be nonlinear in $K_{ij}$, which makes the explicit reversion of velocity $\dot{h}_{ij}$ in terms of the momentum $\pi^{ij}$ impossible. This problem can be solved by rewriting (3) in an equivalent form
@@
where the action $S\left(\phi,N,h_{ij},{}^{3}R_{ij},B_{ij};\mathrm{D}_{i}\right)$ is the same action as (3) with $K_{ij}$ replaced by $B_{ij}$. This strategy of using the Lagrange multiplier to linearize the velocities has been used in [68, 78]. The equations of motion for $\varLambda^{ij}$ ensure the auxiliary field $B_{ij}$ to be exactly identical to $K_{ij}$. Thus the two actions are equivalent, at least at the classical level. By making use of the equation of motion for the auxiliary field $B_{ij}$, we can fix the Lagrangian multiplier $\varLambda^{ij}$ as
@@
Finally, by plugging the solution of $\varLambda^{ij}$, the action (4) can be recast in the form
@@ III.1 Hamiltonian and primary constraints
and the set of their conjugate momenta is denoted by
@@
Since all the velocities $\dot{\varPhi}^{I}$ in the Lagrangian (6) are linearized, the conjugate momenta autonatically correspond to constraints among phase space variables. Indeed, the conjugate momenta are
@@
which result in the primary constraints
@@
Here and throughout this work $\approx$” represents “weak equality”, i.e., an equality that holds only when the constraints are satisfied. On the other hand, the strong equality “==” holds no matter the constraints are satisfied or not.
@@
and the subspace of the phase space when the primary constraints are satisfied as $\varGamma_{\mathrm{P}}$, which can be viewed as a hypersurface in the original phase space $\left\{\varPhi^{I},\varPi_{I}\right\}$.
The canonical Hamiltonian is obtained by performing the Legendre transformation
@@
$=$ $\int\mathrm{d}^{3}x\left(\pi_{i}\pounds_{\vec{\xi}}N^{i}+p\pounds_{\vec{\xi}}\phi+p^{ij}\pounds_{\vec{\xi}}B_{ij}+\pi\pounds_{\vec{\xi}}N+\pi^{ij}\pounds_{\vec{\xi}}h_{ij}\right),$
In the following we briefly mention the properties of $X[\vec{\xi}]$ following [68]. Supposing $Q\left(\vec{x}\right)$ is a general spatial tensor field or density (with spatial indices suppressed for briefness) built of phase space variables $\left\{\varPhi^{I},\varPi_{I}\right\}$, and $\mathcal{F}$ is a scalar functional defined by
@@
with $f\left(\vec{x}\right)$ is a smoothing function. Note $f(\vec{x})$ also constains appropriate spatial indices (suppressed as well) such that $\mathcal{F}$ is invariant under time-independent spatial diffeomorphism. The Poisson bracket of $X[\vec{\xi}]$ and $\mathcal{F}$ are
@@
In (18), the Poisson bracket $\mathcal{F},\mathcal{G}]$ is defined by
$\mathcal{F},\mathcal{G}]\coloneqq\int\mathrm{d}^{3}z\left(\frac{\delta\mathcal{F}}{\delta\varPhi_{I}(\vec{z})}\frac{\delta\mathcal{G}}{\delta\varPi^{I}(\vec{z})}-\frac{\delta\mathcal{F}}{\delta\varPi^{I}(\vec{z})}\frac{\delta\mathcal{G}}{\delta\varPhi_{I}(\vec{z})}\right).$ (19)
The key point is that the right side of the second line of the equation vanishes on the constraint surface if $Q(\vec{x})$ is a constraint.
Through integrations by parts, $X[\vec{\xi}]$ can be written in a more convenient form
@@
As we shall see later, $\mathcal{C}_{i}\approx 0$ are actually the secondary constraints associated with the primary constraints $\pi_{i}\approx 0$. Moreover, the Poisson bracket of $\mathcal{C}_{i}$ with any constraint $Q$ can be obtained by
@@
Due to the presence of primary constraints, the time evolution is determined by the total Hamiltonian defined by
@@
where $\lambda_{I}:=\left\{v^{i},v,v_{ij},\lambda,\lambda_{ij}\right\}$ are the undetermined Lagrange multipliers. Since the canonical Hamiltonian $H_{\mathrm{C}}$ is well-defined only on the subspace defined by the primary constraints $\varGamma_{\mathrm{P}}$, we can directly use $\left.H_{\mathrm{C}}\right|_{\varGamma_{\mathrm{P}}}$ given in (14) instead of $H_{\mathrm{C}}$ in (13) in the subsequent calculations.
@@ III.2 Consistency condition
Constraints must be preserved in time evolution. For the primary constraints, we must require that
@@
which are the so-called consistency conditions for the primary constraints. These conditions may yields further constraints.
With no further requirement on the structure of the Lagrangian in (6), the Poisson brackets between primary constraints have been calculated in detail in [85], which are summarized in Appendix B. For later convenience, we write (25) in matrix form
@@
where $\mathcal{C}_{i}$ is defined in (22). Note $\mathcal{C}_{i}$ must have vanishing Poisson brackets with any other constraints due to the property (22). In the fourth line of (26), $\mathcal{C}$ is defined by
@@
According to (129), $\left[p^{ij}(\vec{x}),\tilde{\pi}^{kl}(\vec{y})\right]$ in the third line in (26) is proportional to the kinetic matrix
@@
which we assume to be non-degenerate. This is because $B_{ij}$ is the auxiliary field corresponding to $K_{ij}$, The non-degeneracy of (29) implies that the kinetic term for $\dot{h}_{ij}$ is not degenerate, which guarantees the theory includes General Relativity as its limit, and in particular, the existence of gravitational waves.
As a result, the Lagrange multiplies $\lambda_{ij}$ can be fixed from the third line in (26) to be
@@
Then the second and the fourth line of (26) yield another two secondary constraints
@@
Constraints $\frac{\delta S}{\delta\phi(\vec{x})}$ and $\mathcal{C}(\vec{x})$ do not generate additional constraints due to consistency conditions, which will be clear in the next section.
After finding all the primary and secondary constraints, we are now ready to classify all the constraints. According to the terminology of Dirac-Bergmann algorithm, a constraint that has vanishing Poisson brackets with all contraints (including itself) is dubbed a first-class constraint, otherwise a second-class constraint. The difference is that each first-class constraint eliminates a single degree of freedom, while a second-class constraint, which reduce one phase space dimension, eliminate half degree of freedom.
According to the above discussion on $\mathcal{C}_{i}$ and constraint algebra in Appendix B, the Poisson brackets among all the constraints can be summarized in the so-called Dirac matrix:
@@
where a “XX” stands for generally non-vanishing Poisson brackets. In total there are 22 constraints which can be divided into two classes:
@@
The number of DOF is thus
$\#_{\mathrm{DOF}}$ $=$ $\frac{1}{2}\left(2\times\#_{\mathrm{var}}-2\times\#_{1\mathrm{st}}-\#_{2\mathrm{nd}}\right)$ (33)
@@
$=$ $3,$
which explicit shows that spatially covariant gravity with temporal derivative arises only through the extrinsic curvature, has 3 propagating degrees of freedom, no matter with or without extra auxiliary field(s).